OpenStudy (anonymous):

mx''+kx=Acos(pt) Is there a general solution to this, not just where there is resonance or similar?

OpenStudy (turingtest):

I don't think I get your terminology, but the form of the answer will depend on the sign of k, which is I assume your damper/amplifier

OpenStudy (anonymous):

Forced oscillator.

OpenStudy (turingtest):

like all second-order non-homogeneous DE's it will have the form\[y=y_c+y_p=c_1y_1+c_2y_2+y_P\]and \(y_c\) depends on the value of k being positive, negative, or zero

OpenStudy (anonymous):

Or, to go for the whole hog, is there an exact solution for\[m\frac{d^2x}{dt^2}=-kx-c\frac{dx}{dt}+acos(\theta t)\]

OpenStudy (anonymous):

@experimentX , do you have any ideas how to tackle the second DE? Despite both of your help, I'm still very much guessing solutions to DE based on aforeknown schemes.

OpenStudy (experimentx):

this is your homogeneous DE. \[ m\frac{d^2x}{dt^2} + c\frac{dx}{dt} + kx = 0\] do you know how to find it's solution?

OpenStudy (anonymous):

Assume \[x=Ae^{pt}\]Divide by\[Ae^{pt}\]Solve the quadratic for p and find A by initial conditions.

OpenStudy (experimentx):

yep ... it can be modelled as harmonic oscillator c/m as your damping constant and sqrt(k/m) as your oscillating frequency

OpenStudy (experimentx):

\[ m\frac{d^2x}{dt^2} + c\frac{dx}{dt} + kx = f(t) \] and f(t) is your input to the system. the solution would be sum of complimentary function (which is the solution of homogeneous DE) and Particular Integral. we add complimentary function because you can test it in DE. i doesn't affect the result.

OpenStudy (experimentx):

And also I'll give you one basic intuition behind \( A e^{pt } \). you can consider is as linear operator on x \[ L(D) x = (D-a)(D-b)y = 0 --(1) \\ \] Let \[ (D- b)y = u ---(2)\] then we have \[ (D-a)u = 0 --- (3)\] this is linear in u ... solve for u, then put it in (2) and solve whole DE. probably you willget in terms of e^(px) is is characteristic of first order DE.

OpenStudy (experimentx):

solving for particular integral is not quite easy. there are bunch of techniques. most used are 1. undetermined coefficients (this is purely guessing) 2. reduction of orders (i like this one) 3. variation of parameters 4. and there is one commonly used ones here (in my place) ...perhaps you guys are not familiar with it.

OpenStudy (anonymous):

Before I read your last & penultimate posts: how do we know that adding particular and complementary is enough? Couldn't there be a huge number of complimentary solutions 'hiding'?

OpenStudy (experimentx):

no there cannot be more that two integral constants. that is from the definition of DE. DE are made by eliminating these constants.

OpenStudy (experimentx):

if you have simple input type of polynomials or exponential or trigonometric ... consider yourself lucky. polynomials >> you can easily guess solution exponential input >> http://ocw.mit.edu/courses/mathematics/18-03sc-differential-equations-fall-2011/unit-ii-second-order-constant-coefficient-linear-equations/exponential-response/ this is easy to remember trigonometric >> can be changed into exponential using Euler's formula. just look for imaginary part (if you have sine) or real part ... if you have cosine.

OpenStudy (anonymous):

What are the commonly used ones for you?

OpenStudy (anonymous):

And thanks for the explanation and link, twas verily useful.

OpenStudy (experimentx):

i don't understand how they do it ... but's it's like using \[ {1 \over D} x = \int x dx\] I'll show you for particular problem \[ (D - 1)(D - 2)x = t \\ P.I. = {1 \over (D - 1)(D - 2)} \;t = {1 \over -2( D - 1)} (1 + {D \over 2} + {D^4 \over 4} + ..)t =\\ -{1 \over 2 (D -1)}(t + 1/2) = {1 \over 2} ( 1 +D + D^2 + ..)(t+ 1/2) = 1/2(t + 1/2) + 1/2\] something like this ... if you encounter 1/D ... you integrate.

OpenStudy (turingtest):

in English this is the annihilator method

OpenStudy (experimentx):

hmm ... really? i didn't know that. Man I'm really bad at English.

OpenStudy (anonymous):

Thanks. And sorry- I'm not familiar with the D notation. D is d/dx?

OpenStudy (experimentx):

yep ... this acts as linear operator.

OpenStudy (turingtest):

no you're not, it's just a piece of terminology If we are in fact talking about the same thing D is the "differential operator" and acts like taking a derivative.

OpenStudy (turingtest):

I'm not so savvy on the technique myself

OpenStudy (anonymous):

Dy =d(y)/dx L(D)y= L(d(y)/dx)?

OpenStudy (anonymous):

I only learnt about operators the other day, so pardon my ineptitude.

OpenStudy (experimentx):

this method gives quick and dirty solution.

OpenStudy (experimentx):

L(D) = L(d/dx) ... not dy/dx ... just d/dx .... this will operate on 'y'

OpenStudy (experimentx):

L(d/dx) is just some function of d/dx

OpenStudy (anonymous):

So functions are written as F(y) , and operators as Fy?

OpenStudy (experimentx):

No ... L(d/dx) whole is operator \( {d^2y \over dx^2} + 2{dy \over dx} + y = ({d^2 \over dx^2} +2 {d \over dx} + 1) y = (D^2 + 2D + 1)y = F(D) y\)

OpenStudy (turingtest):

you choose for instance the operator\[(D^2+4)\]to annihilate\[\sin(2x)\] because\[(D^2+4)\sin(2x)=D^2(\sin(2x))+4\sin(2x)\]\[=D(2\cos(2x))+4\sin(2x)=-4\sin(2x)+4\sin(2x)=0\]hence the name

OpenStudy (experimentx):

I wrote L in-place of F ... because it is linear.

OpenStudy (anonymous):

\[\frac{1}{D}x=\int\limits x dx\] So\[\frac{1}{F}x=y \rightarrow Fy=x\] So is an inverse operator exactly the same as an inverse function? Does the 1/D notation imply that you can algebraically manipulate an equation with D as if it were a constant?

OpenStudy (experimentx):

well ... sure you can manipulate D's to your ease. but careful when dealing with variables. For example in above ... you see i manipulated it as geometric series and got the particular integral. http://www.wolframalpha.com/input/?i=y%27%27+-+3+y%27+%2B+2y+%3D+x similarly you can try this \[ D(D-1)y = x\]

OpenStudy (anonymous):

\[D(D-1)y=\frac{dy}{dx}(\frac{dy}{dx}-1)?\]

OpenStudy (anonymous):

Is this a correct reading of the notation?

OpenStudy (experimentx):

Oops sorry no ... \[ D(D-1)y=(\frac{d}{dx}(\frac{d}{dx}-1) )y \]

OpenStudy (anonymous):

What does that MEAN, though? What do you do to y, and in what order?

OpenStudy (experimentx):

Damn ... chrome just crashed ... therefor no latex.